OpenStudy (anonymous):

What is the equation of the line that passes through the points (4/5,1/5) AND (1/2 ,3/2) ?

OpenStudy (anonymous):

a ) Y= -X -1 B) Y= -X + 1 C) Y = -13/3x + 11/3 d) y= 7/3x + 1/3

OpenStudy (anonymous):

@Luigi0210 @okrzesik10 @phi @Yacoub1993 @greenlegodude57 @francoisderozay @e.mccormick @Vallarylynn16 @DullJackel09 @surjithayer @Ashleyisakitty @mlb004 @okrzesik10 @bookworm00981 @Vallarylynn16 @Callisto @DullJackel09 @greenlegodude57 @hba

OpenStudy (traceur55):

You know this is History right?

OpenStudy (anonymous):

https://www.desmos.com/calculator plug in your answer shoices and see which one matches the question c: